Circular No. 8943

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 J4 (McNAUGHT) R. H. McNaught reports his discovery of yet another comet on CCD

images obtained with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation tabulated below), when the object appeared very diffuse with a coma elongated in p.a. 240^{o} and a diffuse 5′ fan tail spanning 20^{o}-30^{o} that curved to p.a. ∼ 250^{o}. Stacked images from May 11.76 show that a condensation in the very diffuse coma was elongated along p.a. 240^{o} for ∼ 20′′, with a 2′.2 tail in p.a. 230^{o} possibly curving to the south; stacked images in better seeing on May 11.81 showed the head as a diffuse oval patch with the leading edge appearing as almost a point. Additional stacked images from May 12.76 reveal a very diffuse head of size 30′′ × 40′′ with a 2′.0 tail in p.a. 230^{o}; the leading edge of the comet is angled, giving an appearance of just a tail with the brightest part of the comet behind where the nuclear condensation would normally be expected. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, J. C. Pelle reports that he measured a tail 7′ long in p.a. 240^{o} from images taken by N. Teamo (Punaauia, French Polynesia, 0.41-m f/8.0 reflector) on May 12.6; images taken on May 13.6 showed no visible condensation, making for difficult measurement.

2008 UT  α_{2000}  δ_{2000}  Mag.
May 10.79355  23^{h}26^{m}34.^{s}26  −7^{o}36′56′′.3  16.5

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-J69.

T = 2008 June 19.441 TT  ω = 93.^{o}709
Ω = 289.109 }2000.0
q = 0.44485 AU  i = 88.396

COMET 198P/ODAS Comet P/2006 B7 = 1998 X1 (cf. IAUC 8929, 8930, 8935) has been

assigned the number 198P (cf. MPC 62519).

COMET C/2008 J3 (McNAUGHT) Corrigendum. The information on IAUC 8942 from Pelle was concern-

ing comet C/2008 J4 (see above), not C/2008 J3.

2008 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green